On January 1, 2021, Jay Company acquired all the outstanding ownership shares of Zee Company. In assessing Zee’s acquisition-date fair values, Jay concluded that the carrying value of Zee’s long-term debt (8-year remaining life) was less than its fair value by $16,600. At December 31, 2021, Zee Company’s accounts show interest expense of $13,950 and long-term debt of $310,000. What amounts of interest expense and long-term debt should appear on the December 31, 2021, consolidated financial statements of Jay and its subsidiary Zee?

Interest expense Long-term debt
a. $16,025 $326,600
b. $16,025 $324,525
c. $11,875 $326,600
d. $11,875 $324,525

Multiple Choice

  • Option D

  • Option A

  • Option C

  • Option B
Answer #1
Annual Amortization of Excess Debt $ 2,075 =16600/8
Interest Expense on Consolidated financial statements $ 16,025 =13950+2075
Long-term debt on Consolidated financial statements $ 324,525 =310000+16600-2075
Correct answer is option B .